# Learning How to Dynamically Decouple

### v1

@@ III.2.1 Suppressing noise during mid-circuit measurements
Figure 3: Suppressing noise during MCMs through DD. A Bell state is prepared on qubit $q_{0}$ and $q_{2}$. The magenta block on qubit $q_{1}$ is a MCM that is repeated $r$ times. The purple blocks on qubit $q_{0}$ and $q_{2}$ indicate DD sequences inserted during the MCMs to mitigate noise on $q_{0}$ and $q_{2}$. We vary $r$ and study the performance of different DD sequences on a $27$ qubit IBM Quantum Falcon device ($ibm\_hanoi$). The corresponding experimental results are shown in Fig. 4.
@@
As depicted in Fig. 3, we consider the task of preparing the Bell state $|\Phi_{+}\rangle$ between qubit $q_{0}$ and qubit $q_{2}$ while qubit $q_{1}$ is subject to $r$ repeated MCMs. Here, varying $r\in\{1,...,15\}$ allows us to amplify the amount of noise introduced in the quantum circuit.
@@
Figure 4: DD noise suppression during MCMs. Fidelity for preparing a Bell state between qubit $q_{0}$ and $q_{2}$ in Fig. 3 as a function of the number of MCMs performed on qubit $q_{1}$. The dashed red curve shows the fidelity with MCMs but without DD and the dash-dotted purple curve shows the fidelity with a delay equivalent to MCMs. DD sequences are inserted on $q_{0}$ and $q_{2}$ when the qubits are idling. The dashed green curve corresponds to the CPMG sequence, the dash-dotted orange curve corresponds to the XY4 sequence and the solid blue curve corresponds to the optimized LDD sequence. The optimal LDD parameters are shown in Table 1. For comparison, the solid black curve shows the $T_{1}$ decay. The error bars show the interquartile range.
@@
Table 1: Optimal LDD parameters to suppress noise during MCMs. The angles in the LDD gate given in Eq. (9) are optimized to find the optimal rotational (Euler) angles $\vec{x}^{*}=(\theta^{*},\phi^{*},\lambda^{*})$. The optimal angles are shown for a different number of MCMs applied on $q_{1}$ that correspond to the idling times of qubit $q_{0}$ and $q_{2}$. For reference, the Euler angles for the Pauli $X$, and Pauli $Y$ gates used in the XY4 sequence are $(\pi,-0.5\pi,0.5\pi)$ and $(\pi,0,0)$, respectively.
First, we compute $F$ without any DD, once with $r$ MCMs and once with a delay equivalent to $r$ MCMs. The experimental result are shown in Fig. 4 (red and purple curves) where we report the median and the lower and upper quartiles of ten measurements. In the presence of MCMs we observe a large drop in $F$ from $\smash{0.916\begin{subarray}{c}+0.004\\ -0.026\end{subarray}}$ at $r=1$ to ${0.447\begin{subarray}{c}+0.052\\ -0.178\end{subarray}}$ at $r=15$. Without MCMs $F$ drops from ${0.905\begin{subarray}{c}+0.015\\ -0.075\end{subarray}}$ at $r=1$ to ${0.386\begin{subarray}{c}+0.033\\ -0.016\end{subarray}}$ at $r=15$. This implies that a MCM on $q_{1}$ does not have a large impact on $q_{0}$ and $q_{2}$ aside from adding a $820~{}{\rm ns}$ delay. Crucially, the fidelity decrease due to the long delays induced by $r$ MCMs on $q_{1}$ can be mitigated by DD sequences inserted on neighbouring qubits $q_{0}$ and $q_{2}$ during the MCM measurement, see Fig. 3. In Fig. 4 we compare the performance of LDD (blue), CPMG (green), and XY4 (orange). The LDD sequence, whose corresponding optimal parameters are shown in Table 1, yields the best performance, resulting in e.g., a fidelity of ${0.853\begin{subarray}{c}+0.012\\ -0.007\end{subarray}}$ at $r=15$ while CPMG and XY4 result in ${0.814\begin{subarray}{c}+0.012\\ -0.005\end{subarray}}$ and ${0.774\begin{subarray}{c}+0.007\\ -0.021\end{subarray}}$, respectively. To evaluate the reliability of the LDD sequence we compute $F(\vec{x}^{*})$ ten times after the learning and report the lower quartile, median, and upper quartile of these ten runs. Due to queuing, these circuits were executed on the hardware two days after the optimal parameters $\vec{x}^{*}$ were learnt. This indicates that the learned parameters $\vec{x}^{*}$ are stable in time. We attribute the residual decay to $T_{1}$ shown for comparison in Fig. 4 as a solid black curve. This curve is computed as $e^{-t/T_{1}}$ where $t$ is the idling time and $T_{1}=167~{}\mu{\rm s}$ is the average $T_{1}$ time of qubits $q_{0}$ and $q_{2}$ of ibm_hanoi.
III.2.2 Suppressing deep circuit noise
Figure 5: Suppressing deep circuit noise through DD. (a) Qubit coupling graph of the $27$ qubit IBM Falcon device $ibm\_cairo$. Qubits are shown as circles labeled by numbers. The green circles highlight the qubits in the quantum circuit in (b). The grey dashed box shows the chain of qubits $0,1,4,7,10,12,15,18,21$, and $23$ used in the experiment. (b) Quantum circuit to prepare a Bell state between qubit $q_{0}$ (one edge of the chain) and qubit $q_{i}$ (here shown for $i=12$) using a ladder of CNOT gates. The insertion of DD sequences is shown as purple boxes. (c) Fidelity of preparing a Bell state between qubit $q_{0}$ and $q_{i}$ in the chain in the coupling graph (a) as a function of the number of intermediate qubits. The dashed red curve shows the fidelity without DD, the dashed-dotted orange curve corresponds to the XY4 sequence, the dashed-dotted green curve corresponds to the CPMG sequence, and the solid blue curve correspond to the optimized LDD sequence. The optimal LDD parameters are shown in Table 2. The error bars show the interquartile range.
The corresponding coupling graph of the IBM Quantum device is shown in Fig. 5(a) where the considered qubit chain is highlighted in the dashed grey box. In Fig. 5(b) we show the quantum circuit that prepares a Bell state between qubits $q_{0}$ and $q_{12}$ by bringing them into proximity with a ladder of SWAP gates. Since the ancilla qubits are in their ground state |$\ket{0}$, we can implement each SWAP gate with two CNOT gates instead of three. Similar gate ladders with a single CNOT at each rung often occur in quantum simulation algorithms to create unitaries generated by Pauli strings [61, 62]. In Fig. 5(c) we plot the fidelity $F$ for preparing the Bell state between the two edge qubits of the chain in Fig. 5(a) as a function of the intermediate qubits (IQ) shown by the dashed red curve. We see a fidelity decrease from ${0.948\begin{subarray}{c}+0.002\\ -0.002\end{subarray}}$ for a chain consisting of $2$ qubits (i.e., 00 IQ) to ${0.720\begin{subarray}{c}+0.024\\ -0.013\end{subarray}}$ for a chain consisting of $10$ qubits (i.e., $8$ IQ).
@@ III.2.2 Suppressing deep circuit noise
To mitigate the Bell state fidelity decrease, as depicted in Fig. 5(b), we insert DD sequences on idling qubits and compare their performance. The results shown in Fig. 5(c) for the CPMG sequence (green), the XY4 sequence (orange), and the LDD sequence (blue) suggest that the best performance is obtained for LDD where the corresponding optimal parameters are given in Table 2.
In fact, since the fidelity obtained through inserting XY4 and CPMG can even be below the fidelity obtained without DD, observing a significant fidelity drop for XY4 at $1$ IQ, we conclude that DD can increase the noise instead of suppressing it. This situation is avoided by LDD as the DD sequence is tailored to the device.
@@
Table 2: Optimal LDD parameters for suppressing deep circuit noise. The angles in the LDD gate given in Eq. (9) are optimized to find the optimal rotational (Euler) angles $(\theta^{*},\phi^{*},\lambda^{*})$. The optimal $\vec{x}^{*}=(\theta^{*},\phi^{*},\lambda^{*})$ are shown for different number of intermediate qubits (IQ) in the chain configuration in Fig. 5(a).
IV Conclusion
Dynamical decoupling (DD) is a powerful noise suppression strategy that averages out detrimental processes by applying properly designed pulses to the system. We introduced the framework of “learning dynamical decoupling” (LDD). Instead of considering a DD sequence with fixed rotational gates, LDD optimizes directly on quantum hardware the rotational parameters in the DD gates. We compared the performance of such optimized DD sequences with the known DD sequences CPMG and XY4 on IBM Quantum hardware. We found that LDD outperforms both sequences in suppressing noise that occurs during mid-circuit measurements and noise that stems from increasing the depth of a quantum circuit.
The LDD sequences that we studied have by design a small number of single-qubit gates and a fixed number of rotational parameters. While we believe that performance can be increased even further by adding more optimization parameters, the results shown in Fig. 5 for different system sizes (i.e., a different number of intermediate qubits) suggest that the number of LDD parameters can remain constant while achieving a similar performance when the system is scaled. As such, the classical optimization overhead does not need to increase when the system size increases. Therefore, the LDD approach considered here is scalable by design. Furthermore, inserting only a small number of single-qubit gates in LDD or DD on idling qubits to suppress noise is important for current quantum devices. Idle times in quantum circuits typically occur when a subset of all qubits undergo two-qubit gates. Therefore, the shorter the two-qubit gate duration is on a device, the more compact the DD sequence needs to be. To illustrate this consider ibm_torino and ibm_sherbrooke which have a median two-qubit gate duration of $84~{}\mathrm{ns}$ and $533~{}\mathrm{ns}$, respectively  [^1] The numbers are as reported from the backends. Furthermore, ibm_torino implements two-qubit CZ gates with tunable couplers while ibm_sherbrooke uses the cross-resonance interaction.. The duration of both the single-qubit $X$ and $\sqrt{X}$ gates are $32~{}\mathrm{ns}$ and $57~{}\mathrm{ns}$ for ibm_torino and ibm_sherbrooke, respectively. As such, on ibm_torino we can insert up to two $X$ or $Y$ gates or one arbitrary single-qubit rotation during the median two-qubit gate duration. By contrast, on ibm_sherbrooke these numbers become eight and four, respectively. Consequently, as the two-qubit gate duration become comparable with the single-qubit gate duration, short DD sequences become more important.
Acknowledgements.
C. A. acknowledge support from the National Science Foundation (Grant No. 2231328). A. R. and C. A. acknowledge support from Knowledge Enterprise at Arizona State University. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.

### v2

@@ III.2.1 Suppressing noise during mid-circuit measurements
Figure 3: Suppressing noise during MCMs through DD. A Bell state is prepared on qubit $q_{0}$ and $q_{2}$. The magenta block on qubit $q_{1}$ is a MCM that is repeated $r$ times. The purple blocks on qubit $q_{0}$ and $q_{2}$ indicate DD sequences inserted during the MCMs to mitigate noise on $q_{0}$ and $q_{2}$. We vary $r$ and study the performance of different DD sequences on a $127$ qubit Eagle device (ibm_kyiv) and a $27$ qubit Falcon device (ibm_hanoi) reported in Appendix  B. The corresponding experimental results are shown in Fig. 4 and Fig. 7(a).
@@
As depicted in Fig. 3, we consider the task of preparing the Bell state $|\Phi_{+}\rangle$ between qubit $q_{0}$ and qubit $q_{2}$ while qubit $q_{1}$ is subject to $r$ repeated MCMs. Here, varying $r\in\{1,...,15\}$ allows us to amplify the amount of noise introduced in the quantum circuit. To study the effect of DD during noisy measurements we first identify a measurement that introduces noise on its neighboring qubits, see Appendix A, and thus map $(q_{0},q_{1},q_{2})\to(q_{120},q_{121},q_{122})$.
@@
Figure 4: DD noise suppression during MCMs. Fidelity of a Bell state between two qubits as a function of the number of MCMs performed on another qubit, see Fig. 3. The dashed red curve shows the fidelity with MCMs but without DD and the dash-dotted purple curve shows the fidelity with a delay equivalent to MCMs. DD sequences are inserted when the qubits are idling. The dashed green curve corresponds to the CPMG sequence, the dash-dotted orange curve corresponds to the XY4 sequence, the dashed magenta curve corresponds to the UR6 sequence, and the solid blue curve corresponds to the optimized LDD sequence. For comparison, the solid black curve, the solid grey curve, and the dashed grey curve show the effect of $T_{1}$ decay, the $T_{2}$ decay, and the decay due to SPAM errors, respectively. The error bars show the interquartile range.
First, we compute $F$ without any DD, once with $r$ MCMs and once with a delay equivalent to $r$ MCMs. The experimental result are shown in Fig. 4 (red and purple curves) where we report the median and the lower and upper quartiles of ten measurements. In the presence of MCMs we observe a large drop in $F$ at $r=5$ from $\smash{0.813\begin{subarray}{c}+0.016\\ -0.01\end{subarray}}$ to $\smash{0.396\begin{subarray}{c}+0.061\\ -0.096\end{subarray}}$. This implies that a MCM on $q_{121}$ strongly impacts $q_{120}$ and $q_{122}$ aside from adding a $1244~{}{\rm ns}$ delay. Crucially, the fidelity decrease due to the long delays and measurement-induced noise of the $r$ MCMs on $q_{121}$ can be mitigated by DD sequences inserted on neighbouring qubits $q_{120}$ and $q_{122}$ during the MCM measurement, see Fig. 3. In Fig. 4 we compare the performance of LDD (blue), CPMG (green), XY4 (orange), and UR6 (magenta). The LDD sequence yields the best performance, resulting in e.g., a fidelity of $0.631\begin{subarray}{c}+0.007\\ -0.007\end{subarray}$ at $r=15$ while CPMG, XY4, and UR6 result in $0.625\begin{subarray}{c}+0.019\\ -0.045\end{subarray}$, $0.499\begin{subarray}{c}+0.010\\ -0.021\end{subarray}$, and $0.521\begin{subarray}{c}+0.08\\ -0.054\end{subarray}$ respectively. To evaluate the reliability of the LDD sequence we compute $F(\vec{x}^{*})$ ten times after the learning and report the lower quartile, median, and upper quartile of these ten runs. Due to queuing, these circuits were executed on the hardware two days after the optimal parameters $\vec{x}^{*}$ were learnt. This indicates that the learned parameters $\vec{x}^{*}$ are stable in time. We attribute the residual decay in part to $T_{1}$, $T_{2}$, and state preparation and measurement (SPAM) errors shown for comparison in Fig. 4 as a solid black, a solid grey, and a dashed grey curve, respectively. The $T_{1}$ curve is computed as $e^{-t/T_{1}}$ where $t$ is the idling time and $T_{1}=245~{}\mu{\rm s}$ is the average $T_{1}$ time of qubits $q_{120}$ and $q_{122}$ of ibm_kyiv. Similarly, the $T_{2}$ curve is computed as $e^{-t/T_{2}}$ where $T_{2}=175~{}\mu{\rm s}$ is the average $T_{2}$ time of qubits $q_{1}20$ and $q_{1}22$ of ibm_kyiv. The dashed grey line is $0.987e^{-t/T_{2}}e^{-t/T_{1}}$ where $0.987$ is the product of the readout assignment fidelity of qubits $q_{120}$ and $q_{122}$ as reported by the backend.
III.2.2 Suppressing deep circuit noise
Figure 5: Suppressing deep circuit noise through DD. (a) A part of the qubit coupling graph of the $127$ qubit IBM Eagle device ibm_strasbourg. Qubits are shown as circles labeled by numbers. The green circles highlight the qubits in the quantum circuit in (b). The grey dashed box shows the chain of qubits $0,1,2,3,4,5,6,7,8$, and $9$ used in the experiment. (b) Quantum circuit to prepare a Bell state between qubit $q_{0}$ (one edge of the chain) and qubit $q_{i}$ (here shown for $i=5$) using a ladder of CNOT gates. The insertion of DD sequences is shown as purple boxes. (c) Fidelity of preparing a Bell state between qubit $q_{0}$ and $q_{i}$ in the chain in the coupling graph (a) as a function of the number of intermediate qubits. The dashed red curve shows the fidelity without DD, the dashed-dotted orange curve corresponds to the XY4 sequence, the dashed-dotted green curve corresponds to the CPMG sequence, the dashed magenta curve corresponds to the UR6 sequence, and the solid blue curve correspond to the optimized LDD sequence. The error bars show the interquartile range.
The corresponding coupling graph of the IBM Quantum device is shown in Fig. 5(a) where the considered qubit chain is highlighted in the dashed grey box. In Fig. 5(b) we show the quantum circuit that prepares a Bell state between qubits $q_{0}$ and $q_{5}$ by bringing them into proximity with a ladder of SWAP gates. Since the ancilla qubits are in their ground state |$\ket{0}$, we can implement each SWAP gate with two CNOT gates instead of three. Similar gate ladders with a single CNOT at each rung often occur in quantum simulation algorithms to create unitaries generated by Pauli strings [65, 66]. In Fig. 5(c) we plot the fidelity $F$ for preparing the Bell state between the two edge qubits of the chain in Fig. 5(a) as a function of the intermediate qubits (IQ) shown by the dashed red curve. We see a fidelity decrease from first for a chain consisting of $2$ qubits (i.e., 00 IQ) to second for a chain consisting of $10$ qubits (i.e., $8$ IQ).
To mitigate the Bell state fidelity decrease, as depicted in Fig. 5(b), we insert DD sequences on idling qubits and compare their performance. The results shown in Fig. 5(c) for the CPMG sequence (green), the XY4 sequence (orange), the UR6 sequence (magenta), and the LDD sequence (blue) suggest that the best performance is obtained for LDD.
In fact, since the fidelity obtained through inserting XY4 and CPMG can even be below the fidelity obtained without DD, we conclude that DD can increase the noise instead of suppressing it. This situation is avoided by LDD as the DD sequence is tailored to the device.
IV Conclusion
Dynamical decoupling (DD) is a powerful noise suppression strategy that averages out detrimental processes by applying properly designed pulses to the system. We introduced the framework of “learning dynamical decoupling” (LDD). Instead of considering a DD sequence with fixed rotational gates, LDD optimizes directly on quantum hardware the rotational parameters in the DD gates. We compared the performance of such optimized DD sequences with the known DD sequences CPMG, XY4, and UR6 on IBM Quantum hardware. We found that LDD outperforms all three sequences in suppressing noise that occurs during mid-circuit measurements and noise that stems from increasing the depth of a quantum circuit. We believe that this is because the optimal angles in LDD are tailored to the noise of the device. We expect that this performance can be further improved by also optimizing over the DD gate spacings and the number of parameterized gates used. Furthermore, the spacing between the DD gates and the way in which they are compiled into $R_{Z}$, $X$, and $\sqrt{X}$ gates may also impact performance [67].
The LDD sequences that we studied have by design a small number of single-qubit gates and a fixed number of rotational parameters. Appendix C shows that the optimized LDD parameters are robust against perturbations. While we believe that performance can be increased even further by adding more optimization parameters, the results shown in Fig. 5 for different system sizes (i.e., a different number of intermediate qubits) suggest that the number of LDD parameters can remain constant while achieving a similar performance when the system is scaled. As such, the classical optimization overhead does not need to increase when the system size increases. Therefore, the LDD approach considered here is scalable by design. Furthermore, inserting only a small number of single-qubit gates in LDD or DD on idling qubits to suppress noise is important for current quantum devices. Idle times in quantum circuits typically occur when a subset of all qubits undergo two-qubit gates. Therefore, the shorter the two-qubit gate duration is on a device, the more compact the DD sequence needs to be. To illustrate this consider ibm_torino and ibm_sherbrooke which have a median two-qubit gate duration of $84~{}\mathrm{ns}$ and $533~{}\mathrm{ns}$, respectively  [^1] The numbers are as reported from the backends. Furthermore, ibm_torino implements two-qubit CZ gates with tunable couplers while ibm_sherbrooke uses the cross-resonance interaction.. The duration of both the single-qubit $X$ and $\sqrt{X}$ gates are $32~{}\mathrm{ns}$ and $57~{}\mathrm{ns}$ for ibm_torino and ibm_sherbrooke, respectively. As such, on ibm_torino we can insert up to two $X$ or $Y$ gates or one arbitrary single-qubit rotation during the median two-qubit gate duration. By contrast, on ibm_sherbrooke these numbers become eight and four, respectively. Consequently, as the two-qubit gate duration become comparable with the single-qubit gate duration, short DD sequences become more important.
Acknowledgements.
C. A. acknowledges support from the National Science Foundation (Grant No. 2231328). A. R. and C. A. acknowledge support from Knowledge Enterprise at Arizona State University. D. J. E. acknowledges funding within the HPQC project by the Austrian Research Promotion Agency (FFG, project number 897481) supported by the European Union – NextGenerationEU. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or IBM Quantum.
Appendix A Identifying qubits with the most MCM-induced noise
The amount of noise an MCM introduces depends on the qubit on which the MCM is performed. To demonstrate that LDD can suppress noise introduced through MCM’s, we identify qubits in the device where MCM’s have a large effect on the fidelity of a Bell state. This is achieved by identifying qubits on the 127 qubit IBM device (ibm_kyiv) for which repeated MCM’s significantly introduce more noise than letting the qubit idle for the corresponding amount of time. These qubits were found by preparing a Bell state between two qubits, followed by inserting a MCM ($r=1$) or the delay equivalent to a MCM ($r=0$) on another qubit, as shown in Fig. 3. In the experiment, we choose the indices of the three qubit such that their corresponding physical qubits are adjacent to each other. This ensures that the noise in the circuit is introduced by the MCM rather than by the circuit transpilation. The Bell state fidelity is shown as a function of the qubit index on which the MCM or delay are applied in Fig. 6. The indices of the qubits with MCM that we consider in ibm_kyiv are $1,4,7,10,19,22,25,28,31,40,43,46,49,58,61,64,67,76,79,82,85,88,97,100,103,106,115,118,121$ and $124$. These choices ensure that the total three-qubit system forms a chain with nearest neighbour interactions in the coupling map of the device to avoid any additional SWAP gates that may introduce more noise.
The red curve in Fig. 6 shows the state fidelity in the presence of a MCM while the blue curve corresponds to the same circuit without an MCM, i.e., the MCM is replaced by a delay with the same duration. The qubit candidate (index $121$) chosen for the MCM experiments whose results are shown in Fig. 4 in the main text is highlighted by a green ellipse.
Figure 6: Noisy MCM identification experiment. Bell state fidelity between qubits $q_{0}$ and $q_{2}$ as a function of the index of qubit $q_{1}$ in Fig. 3 of the manuscript. The green ellipse shows qubit $121$ for which the state fidelity with a MCM (red curve) is significantly lower than when there is a delay equivalent to a MCM (blue curve).
Appendix B Additional data
B.1 MCM noise
In addition to the data presented in the main text obtained from the 127 qubit Eagle device ibm_kyiv, we also gathered data from the 27 qubit Falcon device ibm_hanoi for LDD, CPMG, and XY4. Here, we prepared Bell state between qubit $q_{0}$ and $q_{2}$, inserted MCM on qubit $q_{1}$. The results for the effect of a MCM on qubit $q_{0}$ and $q_{2}$ are shown in Fig. 7(a). We report the median and the lower and upper quartiles of ten measurements. In the presence of MCMs (red curve), we observe a large drop in $F$ from $\smash{0.916\begin{subarray}{c}+0.004\\ -0.026\end{subarray}}$ at $r=1$ to ${0.521\begin{subarray}{c}+0.052\\ -0.178\end{subarray}}$ at $r=15$. Without MCMs (purple curve) $F$ drops from ${0.905\begin{subarray}{c}+0.015\\ -0.075\end{subarray}}$ at $r=1$ to ${0.374\begin{subarray}{c}+0.033\\ -0.016\end{subarray}}$ at $r=15$. This implies that on ibm_hanoi a MCM on a single qubit does not have a large impact on the adjacent qubits aside from adding a $820~{}{\rm ns}$ delay. These qubits were not selected using the procedure used to identify noisy MCMs on ibm_kyiv.
In Fig. 7(a), we compare the performance of LDD (blue), CPMG (green), and XY4 (orange). The LDD sequence, whose corresponding optimal parameters are shown in Table 1, yields the best performance, resulting in e.g., a fidelity of ${0.855\begin{subarray}{c}+0.012\\ -0.007\end{subarray}}$ at $r=15$ while CPMG and XY4 result in ${0.811\begin{subarray}{c}+0.012\\ -0.005\end{subarray}}$ and ${0.782\begin{subarray}{c}+0.007\\ -0.021\end{subarray}}$, respectively. We attribute the residual decay to $T_{1}$ shown for comparison in Fig. 7(a) as a solid black curve, where $T_{1}=167~{}\mu{\rm s}$ is the average $T_{1}$ time of qubits $q_{0}$ and $q_{2}$ of ibm_hanoi.
Figure 7: DD noise suppression during MCMs and deep circuit noise suppression using ibm_hanoi and ibm_cairo respectively. (a) Fidelity as a function of the number of MCMs performed. The dashed red curve shows the fidelity with MCMs but without DD and the dash-dotted purple curve shows the fidelity with a delay equivalent to MCMs. The dashed green curve corresponds to the CPMG sequence, the dash-dotted orange curve corresponds to the XY4 sequence and the solid blue curve corresponds to the optimized LDD sequence. The optimal LDD parameters are shown in Table 1. For comparison, the solid black curve shows the $T_{1}$ decay. The error bars show the interquartile range. (b) Fidelity of preparing a Bell state between two qubits in the chain in the coupling graph as a function of the number of intermediate qubits. The dashed red curve shows the fidelity without DD, the dashed-dotted orange curve corresponds to the XY4 sequence, the dashed-dotted green curve corresponds to the CPMG sequence, and the solid blue curve correspond to the optimized LDD sequence. The optimal LDD parameters are shown in Table 2. The error bars show the interquartile range.
@@ B.1 MCM noise
Table 1: Optimal LDD parameters to suppress noise during MCMs. The angles in the LDD gate given in Eq. (10) are optimized to find the optimal rotational (Euler) angles $\vec{x}^{*}=(\theta^{*},\phi^{*},\lambda^{*})$. The optimal angles are shown for a different number of MCMs applied on $q_{1}$ that correspond to the idling times of qubit $q_{0}$ and $q_{2}$. For reference, the Euler angles for the Pauli $X$, and Pauli $Y$ gates used in the XY4 sequence are $(\pi,\pi,0)$ and $(\pi,0,0)$, respectively.
B.2 Deep circuit noise
In addition to the data presented in the main text on the 127 qubit Eagle device ibm_strasbourg, we also gathered data on the 27 qubit Falcon device ibm_cairo for LDD, CPMG, and XY4, see Fig. 7(b). We plot the fidelity $F$ for preparing the Bell state between the two edge qubits of the qubit chain in ibm_cairo as a function of the intermediate qubits (IQ) shown by the dashed red curve. We see a fidelity decrease from ${0.948\begin{subarray}{c}+0.002\\ -0.002\end{subarray}}$ for a chain consisting of $2$ qubits (i.e., 00 IQ) to ${0.720\begin{subarray}{c}+0.024\\ -0.013\end{subarray}}$ for a chain consisting of $10$ qubits (i.e., $8$ IQ).
@@ B.2 Deep circuit noise
To mitigate the Bell state fidelity decrease, we insert DD sequences on idling qubits and compare their performance. The results shown in Fig. 7(b) for the CPMG sequence (green), the XY4 sequence (orange), and the LDD sequence (blue) suggest that the best performance is obtained for LDD where the corresponding optimal parameters are given in Table 2.
@@
Table 2: Optimal LDD parameters for suppressing deep circuit noise. The angles in the LDD gate given in Eq. (10) are optimized to find the optimal rotational (Euler) angles $(\theta^{*},\phi^{*},\lambda^{*})$. The optimal $\vec{x}^{*}=(\theta^{*},\phi^{*},\lambda^{*})$ are shown for different number of intermediate qubits (IQ).
Appendix C Robustness of the optimal angles to perturbations
To study the robustness of the optimized angles we perform additional MCM experiments on ibm_kyiv.
Figure 8: Robustness of the optimal angles to pertubations. The Bell state fidelity between qubits $q_{120}$ and $q_{122}$ as a function of the strength of the perturbation $\epsilon$, see Eq. (11) and Eq. (12). The fidelity corresponding to the optimal angles ($\epsilon=0$) is shown as a grey line.
We first optimize the angles $\theta,\phi,\lambda$ in a LDD sequence for $r=1$ as in Fig. 3. Next, we investigate the robustness by perturbing the angles in $R_{\vec{x}}=R(\theta,\phi,\lambda)$ and reporting the effect on the fidelity.
In particular, we uniformly sample three parameters $\Delta\theta$, $\Delta\phi$, and $\Delta\lambda$ from the interval $[-2\pi,2\pi]$. We normalize the resulting vector $\vec{\delta}=[\Delta\theta,\Delta\phi,\Delta\lambda]$ to have a norm of $\epsilon$ which controls the magnitude of the perturbation, i.e.,
$\vec{\Delta}=\epsilon\ \frac{\vec{\delta}}{||\vec{\delta}||}=[\Delta\theta^{{}^{\prime}},\Delta\phi^{{}^{\prime}},\Delta\lambda^{{}^{\prime}}].$ (11)
We perform experiments for the perturbed LDD gates,
$R_{\vec{x}+\vec{\Delta}}=R(\theta+\Delta\theta^{{}^{\prime}},\phi+\Delta\phi^{{}^{\prime}},\lambda+\Delta\lambda^{{}^{\prime}}),$ (12)
as a function of the perturbation strength $\epsilon$. The results are shown in Fig. 8 where each data point corresponds the median of $10$ samples and the error bars show the upper and lower quartile range. For comparison we show the fidelity found without perturbing the optimized angles as a grey line.
Appendix D DD transpiled circuits
The circuits in this work are transpiled to the qubit coupling map and the hardware native gates of the quantum device. To insert a DD sequence in a circuit we first schedule the circuit. I.e., we attach to the hardware native circuit instructions their corresponding duration and schedule them as late as possible. Next, idling times on the qubits that are long enough are replaced by delays and DD gates. The duration of the delays are determined by the total idle time and the DD sequence. In this appendix we exemplify the gate sequences resulting from this transpilation process for a delay of $1244~{}{\rm ns}$ which corresponds to the duration of a mid-circuit measurement on ibm_kyiv, see also Sec. III.2.1. We implement CPMG, XY4, UR6, and LDD to fit this idling time. Fig. 9 shows the transpiled circuits and Fig. 10 shows their timeline diagram with time measured in units of system cycle time.
Figure 9: Transpiled DD sequences. (a) CPMG, (b) XY4, (c) UR6, and (d) LDD inserted in an idle time of the same length as a mid-circuit measurement of duration $1244~{}{\rm ns}$, i.e., $5600[dt]$ where $dt=0.22\ n$s. The hardware native gates are $X,\sqrt{X},\ \text{and}\ R_{Z}(\theta)$.
Figure 10: Timeline drawing of the transpiled DD sequences. The (a) CPMG, (b) XY4, (c) UR6, and (d) LDD sequences are the same as those in Fig. 9 and plotted as a function of time in units of $dt$ on the $x$-axis. Here, the width of the $X$ and $\sqrt{X}$ gates, shown in green and red, respectively, matches their duration. The $R_{Z}(\theta)$ gates, shown as curved arrows, are virtual and thus have a duration of zero.